Dickson, Inc., has a debt-equity ratio of 2.25. The firm’s weighted average cost of capital is 10 percent and its pretax cost of debt is 7 percent. The tax rate is 22 percent.

a. What is the company’s cost of equity capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
b. What is the company’s unlevered cost of equity capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
c. What would the company’s weighted average cost of capital be if the company's debt-equity ratio were .75 and 1.25? (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

WACC = We x ke + Wd x Kd(after tax) 0.10 = (1/3.25) x ke +(2.25/3.25) x (0.0546) Ke = 20.22% Cost of Equity Re = Ra +(Ra-Rd)(Ra = unlevered cost of equity If Debt equity = 1.25 cost of equity = Re - Ra +(Ra-Rd)(D/E)(1-t) = 16.47% 0.118+((0.118-0.07)*